OpenStudy (anonymous):

Given the point (3,4) and the slope of 6, find y when x = 27 What form do I use for this?

OpenStudy (luigi0210):

Point-slope i think..: \[\LARGE y-4=6(3-27)\]

OpenStudy (anonymous):

Which form were you given?

OpenStudy (anonymous):

We can use the Point-Slope form of a line: y−yo=m(x−xo) We were given the slope: m=6 And a starting point, (xo,yo)=(3,4) And an x value, x=27

OpenStudy (anonymous):

I think Luigi0210 is right.

OpenStudy (anonymous):

y = 6x-14 y = 6(27)-14 y = 148 I got something different

OpenStudy (anonymous):

y-4=6(x-3) y-4=6x-18 y=6x-14

OpenStudy (luigi0210):

Whoops, got it backwards: \[\LARGE y-4=6(27-3)\]

OpenStudy (anonymous):

Yes, the answer was 148, thank you so much everyone! I really appreciate it

OpenStudy (anonymous):

Thanks for typing out the equation and explaining it to me, too!

OpenStudy (luigi0210):

Yup~